A flat belt is required to transmit 22 kW from a 250 mm diameter pulley running at 1450 rpm to a 355 mm diameter pulley. The proposed belt is 3.5 mm thick. The other system specifications are as follows: Coefficient of friction 0.7, Density of the belt is 1100 kg/m3, Maximum permissible stress is 7 MPa. The distance between the shafts is 1.8 m.


Calculate the width required. 


1
Expert's answer
2021-09-19T00:20:24-0400

solution:


"V=\frac{\pi dn}{60\times1000}"


"V=\frac{\pi\times250\times1450}{60\times1000}"

"=18.98m\/s=19m\/s"


V"\implies" belt velocity "(m\/s)"


"\alpha_s=180-2sin^{-1}(\frac{D-d}{2C})"


"\alpha_s=180-2sin^{-1}(\frac{0.355-0.25}{2\times1.8})"


"=176.66\degree"

in radians; "\alpha_s=(\frac{176.66}{180})\pi=3.08 radians"


according to the power transmitted by the belt


"P=\frac{(P_1-P_2)V}{1000}"


"22=\frac{(P_1-P_2)19}{1000}"


"P_1-P_2=1157.9(N)"

by solving the upper equation


"P_1=1388.2N , P_2=230.3N"

"b=56.66mm=57mm"